OpenStudy (anonymous):

Which equation describes the four ordered pairs (`1,3), (0,1), (2,9), (3,13)? A. y=2x+1 B. y=4x+1 C. y=4x-1 D. y=2x-1

OpenStudy (anonymous):

Which equation describes the four ordered pairs (-1,-3), (0,1), (2,9), (3,13)? A. y=2x+1 B. y=4x+1 C. y=4x-1 D. y=2x-1